Two point charges lie on the x axis. A charge of +1.5
Answer #1

apply eelectric field E = Kq/r^2   or each charge

so

a. ENet at -2cm =    E1+E2

Enet = ( 9e9 * 1.5 e-6/(0.02*0.02)   +(9e9 * 6.6 e-6/(0.12*0.12)

Enet = 3.787*10^7 N/C

----------------------------

b. ENet at 2cm =    E1+E2

Enet = ( 9e9 * 1.5 e-6/(0.12*0.12)   +(9e9 * 6.6 e-6/(0.02*0.02)

Enet = 1.49 *10^8 N/C